# Scaling LLM Planning: NL2FLOW for Parametric Problem Generation and Rigorous Evaluation

### v3

@@ 2.3 Prompt Generation
Develop a concise plan to achieve the goals. The shorter plan is preferable. A plan consists of individual actions, each represented by its outputs and an action name followed by its parameters in parentheses (e.g., outputs = <action-name>(<parameters>). When it becomes impossible to achieve the objectives, the appropriate course of action should be considered as no plan
@@
Listing 4: Concise planning prompt
@@
Develop a concise plan to achieve the goals. The shorter plan is preferable. A plan consists of individual actions, each represented by its outputs and an action name followed by its parameters in parentheses (e.g., outputs = <action-name>(<parameters>). When it becomes impossible to achieve the objectives, the appropriate course of action should be considered as no plan
@@
This study examined the capacity of LLMs to generate executable plans to achieve specified goals and to translate natural language problem descriptions into JSON schemas. I adopted a few-shot prompting approach, constructing prompt styles for both tasks and providing six example demonstrations. Crucially, each example within the prompt presented a natural language description of the plan execution before the corresponding plan expressed in the target format (Listing 5). This was intended to guide the LLM by demonstrating the logical connection between action sequences and their executable representations, thereby increasing the likelihood of generating valid and goal-directed plans.
Listing 5: Planning response example
PLAN EXPLANATION
PLAN
[0] v__1 = a__2()
[1] v__2 = a__1(v__1)
[2] a__3(v__2)
For the planning task, I investigated the impact of prompt style, comparing ’verbose’ (Listing 3) and ’concise’ (Listing 4) prompt styles. The ’verbose’ prompts (mean length 13,413 characters) included a detailed, natural language description of all relevant elements within the planning problem, aiming to provide the LLM with comprehensive contextual information. Conversely, ’concise’ prompts (mean length 9,674 characters) presented these elements as a concise, enumerated list, minimizing verbose text and focusing on key parameters. Listings 6 and 7 show the description of custom action a__1a\_\_1 in verbose and concise prompts, respectively. This variation was designed to assess the LLM’s sensitivity to the level of contextual detail provided in the prompt.
Listing 6: Custom action in verbose prompt
Listing 7: Custom action in concise prompt
name: a__1
inputs: v__11, v__12, v__13
outputs: v__15, v__16, v__18
For translation, a single prompt style (Listing 8) was used, designed to elicit accurate and fluent translations. The complete prompt styles used for each task – verbose_planning_prompt.txt, concise_planning_prompt.txt, and translation_prompt.txt – are included in the supplementary material to facilitate reproducibility, enable independent verification of my results, and encourage further research in this area. Each prompt style followed a consistent structure comprising three sections: Context, Instruction, and Examples (Listings 3 and 4). The Context section provided a programmatically generated natural language description of the problem. The Instruction section detailed the desired planning process, specifying preferences for plan structure, the required output format, and guidance for handling unsolvable problems. To facilitate parsing and execution, plans were to be expressed in a Python-like syntax (e.g., outputs = ⟨action name⟩(⟨parameters⟩)). I standardized action names as a__⟨number⟩and variable names as v__⟨number⟩to promote consistency and reduce ambiguity.
Listing 8: Translation prompt
@@
2.4 Language Model Response Parsing
@@
For the translation task, model outputs were structured as JSON strings representing a planning problem. These strings were then parsed using PyDantic, a data validation and settings management library, to ensure adherence to a predefined schema. This structured format enabled automated evaluation metrics and facilitated downstream analysis. The initial intermediate representation schema used nested structures to represent custom actions; however, this complexity proved unnecessary, as the same information could be represented with a simplified schema. To evaluate the capacity of LLMs to translate natural language descriptions of workflow problems into JSON, I used a streamlined schema (Listing 9) defined as follows:
  • actionsactions represents custom actions.
@@
Listing 9: JSON schema used for evaluating LLM’s translation capacity
actions”: [
id”: a__1”,
input”: [
v__1”,
v__2
output”: [
v__3
mappings”: [],
available_data”: [
v__5”,
askable_parameters”: [
v__1”,
unaskable_parameters”: [
v__4”,
goal_action_ids”: [
a__3
@@
2.5 AI Planning
@@ 2.6.4 Functions
2.6.5 Actions
The domain defines two action types. Custom actions are named according to the pattern a__xa\_\_x.
@@ 2.6.5 Actions
Listing 10: ’custom’ action example in PDDL
@@
For example, Listing 10 illustrate the structure of ’custom’ actions defined in the domain model. This action, a__2a\_\_2, has no parameters. Its precondition requires that the action hasn’t been performed in the past (not(has_donea__2past)not(has\_done\>a\_\_2\>past)) and that a variable v__7v\_\_7 is known with certainty (knownv__7certainknown\>v\_\_7\>certain). The action’s effect marks the action as done in the present (has_donea__2presenthas\_done\>a\_\_2\>present), indicates v__7v\_\_7 has been used, makes v__9v\_\_9 free, establishes v__9v\_\_9 as known with certainty, indicates v__9v\_\_9 is no longer mapped, and increases the $total\mbox{-}cost$ by 10.
The askask action prompts the system to request information from the user if parameters are slot-fillable (Listing 11).
Listing 11: ’ask’ action example in PDDL
@@ 2.6.6 Problem
  • Initial state: The starting situation for the planner. This is a description of what’s true at the beginning.
@@
In this study, the initial state was defined by a set of variables (v__xv\_\_x), each assigned a slot_goodnessslot\_goodness value of 1000000 if slot-fillable, and 150000 if not. These differing slot_goodnessslot\_goodness values represented a ’soft’ constraint, which was not used in this study because the precondition for the ’ask’ action prevented the system from requesting information for variables that were not slot-fillable. The mapped_tomapped\_to predicate establishes a self-mapping for each resource, essentially stating that each resource is initially assigned to itself. A ‘new_item‘ predicate indicates the creation of a new resource. The goal is to achieve conditions to complete a task or tasks ((has_donea__xpresent)and(has_donea__xpresent)(has\_done\>a\_\_x\>present)and(has\_done\>a\_\_x\>present)).
2.7 Plan Evaluation
NL2Flow evaluated plans generated by LLMs using the Planning Domain Definition Language (PDDL). LLM-generated plans were extracted from text and converted to PDDL via NL2Flow, using an intermediate representation from the data generator. The resulting PDDL formulations were then assessed with a symbolic planner. For clarity, a sound plan is defined as one that is executable; a valid plan is executable and achieves the goal state; and an optimal plan is executable, reaches the goal state, and does so with minimal cost.
To assess the soundness of generated plans, the pddl compilation omits goal achievement. NL2Flow’s approach distinguishes between soundness and validity. For soundness checking, the planning problem was compiled into PDDL, but instead of including the original goals as requirements, the actions reported as executed within the LLM’s response were incorporated. This approach tests the LLM’s claimed actions actually executable. Validity was checked by compiling with both original goals and actions in LLM’s plan as requirements; failure to achieve goals indicates a non-valid plan where a planner finds a longer optimal plan. In this low-difficulty workflow problem setting, a plan was invalid, also, if the system prompted the user for a non-slot-fillable variable, or if a ’custom’ action was invoked with unknown input parameter values. I investigated the use of VAL [5] for plan validation; however, VAL did not support the PDDL language features necessary to represent the complexities introduced by NL2Flow’s PDDL compilation process, precluding its use in this evaluation.

### v4

@@ 5 Conclusion
Develop a concise plan to achieve the goals. The shorter plan is preferable. A plan consists of individual actions, each represented by its outputs and an action name followed by its parameters in parentheses (e.g., outputs = <action-name>(<parameters>). When it becomes impossible to achieve the objectives, the appropriate course of action should be considered as ’noplan’
@@
Listing A6: Concise planning prompt
@@
Develop a concise plan to achieve the goals. The shorter plan is preferable. A plan consists of individual actions, each represented by its outputs and an action name followed by its parameters in parentheses (e.g., outputs = <action-name>(<parameters>). When it becomes impossible to achieve the objectives, the appropriate course of action should be considered as ’noplan’
@@
Listing A7: Translation prompt
@@
Language Model Response Parsing
@@
For the translation task, model outputs were structured as JSON strings representing a planning problem. These strings were then parsed using PyDantic, a data validation and settings management library, to ensure adherence to a predefined schema. This structured format enabled automated evaluation metrics and facilitated downstream analysis. The initial intermediate representation schema used nested structures to represent ‘custom‘ actions; however, this complexity proved unnecessary, as the same information could be represented with a simplified schema. To evaluate the capacity of LLMs to translate natural language descriptions of workflow problems into JSON, I used a streamlined schema (Listing A8) defined as follows:
  • actionsactions represents ‘custom‘ actions.
@@
Listing A8: JSON schema used for evaluating LLM’s translation capacity
"actions": [
"id": "a__1",
"input": [
"v__1",
"v__2"
"output": [
"v__3"
"mappings": [],
"available_data": [
"v__5",
"askable_parameters": [
"v__1",
"unaskable_parameters": [
"v__4",
"goal_action_ids": [
"a__3"
AI Planning
@@
Actions
The domain defines two action types. ‘custom‘ actions are named according to the pattern a__xa\_\_x.
@@
Listing A9: ‘custom‘ action example in PDDL
@@
For example, Listing A9 illustrate the structure of ‘custom‘ actions defined in the domain model. This action, a__2a\_\_2, has no parameters. Its precondition requires that the action hasn’t been performed in the past (not(has_donea__2past)not(has\_done\>a\_\_2\>past)) and that a variable v__7v\_\_7 is known with certainty (knownv__7certainknown\>v\_\_7\>certain). The action’s effect marks the action as done in the present (has_donea__2presenthas\_done\>a\_\_2\>present), indicates v__7v\_\_7 has been used, makes v__9v\_\_9 free, establishes v__9v\_\_9 as known with certainty, indicates v__9v\_\_9 is no longer mapped, and increases the $total\mbox{-}cost$ by 10.
The askask action prompts the system to request information from the user if parameters are slot-fillable (Listing A10).
Listing A10: ‘ask‘ action example in PDDL
@@
  • Initial state: The starting situation for the planner. This is a description of what’s true at the beginning. Predicates not explicitly defined as true in the initial state of a problem are considered false (closed-world assumption).
@@
The initial condition of the problem, as described in ‘problem.pddl‘ in supplementary material, can be summarized as follows:
  • slot_goodnessslot\_goodness is defined for all generic objects (v__xv\_\_x and new_object_generic_0new\_object\_generic\_0).
  • totalcosttotal-cost is initialized to 0.0.
  • For any generic object v__xv\_\_x used as an input parameter to a ‘custom‘ action, the predicate $(been_{u}sed\>v\_\_x)$ is true.
  • The predicate (new_itemnew_object_generic_0)(new\_item\>new\_object\_generic\_0) is true.
  • For all generic object, predicates (mapped_to?x?x)(mapped\_to\>?x\>?x) are true. In other words, All generic objects are self-mapped.
@@
In this study, the initial state was defined by a set of variables (v__xv\_\_x), each assigned a slot_goodnessslot\_goodness value of 1000000 if slot-fillable, and 150000 if not. These differing slot_goodnessslot\_goodness values represented a ’soft’ constraint, which was not used in this study because the precondition for the ‘ask‘ action prevented the system from requesting information for variables that were not slot-fillable. The mapped_tomapped\_to predicate establishes a self-mapping for each resource, essentially stating that each resource is initially assigned to itself. A new_itemnew\_item predicate indicates the creation of a new resource. The goal is to achieve conditions to complete a task or tasks (and(has_donea__xpresent)(has_donea__xpresent)and\>(has\_done\>a\_\_x\>present)(has\_done\>a\_\_x\>present)) while minimizing the totalcosttotal-cost.
Regression Analysis Results
Table A1: Logistic regression analysis for no plan with verbose prompt style
DeepSeek-V3
Pseudo R-squ Log-Likelihood
-0.1845 -48.236
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -20.9808 14.949 -1.403 0.16 -50.281 8.319
arity -19.1153 15.123 -1.264 0.206 -48.755 10.524
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals -2.3799 11.24 -0.212 0.832 -24.409 19.649
Granite-3.3-8B-instruct
Pseudo R-squ Log-Likelihood
-0.0137 -146.9
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -0.0534 10.769 -0.005 0.996 -21.161 21.054
arity -15.7939 11.914 -1.326 0.185 -39.144 7.556
coupling -6.1441 4.01 -1.532 0.126 -14.004 1.716
slot-fillable variables -8.00E-17 11.144 -7.18E-18 1 -21.841 21.841
goals 0 nan nan nan nan nan
Llama-3.1-405B-instruct-fp8
Pseudo R-squ Log-Likelihood
-0.03309 -110.76
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -14.7038 12.09 -1.216 0.224 -38.4 8.993
arity -11.7652 13.046 -0.902 0.367 -37.335 13.805
coupling 0 nan nan nan nan nan
slot-fillable variables -2.3115 13.035 -0.177 0.859 -27.86 23.237
goals 0 nan nan nan nan nan
Llama-3.3-70B-instruct
Pseudo R-squ Log-Likelihood
0.007233 -189.65
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions 11.4596 8.745 1.31 0.19 -5.681 28.6
arity -7.33E-17 8.774 -8.36E-18 1 -17.196 17.196
coupling -2.0446 3.507 -0.583 0.56 -8.918 4.829
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Llama-3.1-8B-instruct
Pseudo R-squ Log-Likelihood
-0.08881 -65.393
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -21.3022 13.77 -1.547 0.122 -48.291 5.686
arity -17.1386 13.92 -1.231 0.218 -44.421 10.144
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals -8.56E-17 10.337 -8.28E-18 1 -20.261 20.261
Mixtral-8x22B-instruct-v0.1
Pseudo R-squ Log-Likelihood
-0.05837 -57.104
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -37.7334 5.594 -6.746 0 -48.697 -26.77
arity 0 nan nan nan nan nan
coupling -3.1613 5.535 -0.571 0.568 -14.009 7.687
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Qwen2.5-72B-instruct
Pseudo R-squ Log-Likelihood
-0.01343 -92.919
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -0.7135 12.438 -0.057 0.954 -25.091 23.664
arity -24.2876 12.676 -1.916 0.055 -49.132 0.557
coupling -6.8198 4.877 -1.398 0.162 -16.378 2.739
slot-fillable variables 0 nan nan nan nan nan
goals -1.7047 9.061 -0.188 0.851 -19.464 16.054
Table A2: Logistic regression analysis for no plan with concise prompt style
DeepSeek-V3
Pseudo R-squ Log-Likelihood
inf -18.971
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -27.0837 17.651 -1.534 0.125 -61.679 7.512
arity -22.9006 17.806 -1.286 0.198 -57.800 11.999
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals -2.53e-16 13.455 -1.88e-17 1.000 -26.372 26.372
Granite-3.3-8B-instruct
Pseudo R-squ Log-Likelihood
0.0352 -99.331
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -4.953 4.509 -1.099 0.272 -13.79 3.884
slot-fillable variables -10.9061 7.956 -1.371 0.17 -26.5 4.688
goals -16.5726 7.71 -2.15 0.032 -31.684 -1.461
Llama-3.1-405B-instruct-fp8
Pseudo R-squ Log-Likelihood
-0.02809 -61.746
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -4.3217 15.636 -0.276 0.782 -34.968 26.325
arity -1.1087 16.917 -0.066 0.948 -34.266 32.049
coupling -1.6357 5.5 -0.297 0.766 -12.416 9.144
slot-fillable variables -15.4646 16.811 -0.92 0.358 -48.414 17.485
goals -17.7946 10.376 -1.715 0.086 -38.131 2.542
Llama-3-3-70B-instruct
Pseudo R-squ Log-Likelihood
0.005184 -194.65
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -4.4632 3.174 -1.406 0.16 -10.684 1.757
slot-fillable variables 0 nan nan nan nan nan
goals -5.4489 3.137 -1.737 0.082 -11.597 0.699
Llama-3.1-8B-instruct
Pseudo R-squ Log-Likelihood
-0.03562 -106.62
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -3.6785 11.656 -0.316 0.752 -26.524 19.167
arity -19.0184 11.842 -1.606 0.108 -42.228 4.191
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals -7.0985 8.556 -0.83 0.407 -23.868 9.671
Mixtral-8x22B-instruct-v0.1
Pseudo R-squ Log-Likelihood
0.09344 -91.359
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -7.5232 4.342 -1.733 0.083 -16.033 0.986
slot-fillable variables 0 nan nan nan nan nan
goals -27.1984 4.446 -6.118 0 -35.912 -18.485
Qwen2.5-72B-instruct
Pseudo R-squ Log-Likelihood
-0.1284 -64.37
coef std err z $P>|z|$ [0.025 0.975]
‘custom‘ actions -18.1791 14.71 -1.236 0.217 -47.01 10.652
arity -14.1435 15.761 -0.897 0.37 -45.034 16.747
coupling 0 nan nan nan nan nan
slot-fillable variables -6.4428 16.611 -0.388 0.698 -39 26.114
goals 0 nan nan nan nan nan
Table A3: Logistic regression analysis for validity of plans for problems with feasible plans with verbose prompt style
Deepseek-v3
Pseudo R-squ Log-Likelihood
0.08473 -1043.2
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -41.7807 5.085 -8.216 0 -51.748 -31.814
‘custom‘ actions 62.4231 6.135 10.174 0 50.398 74.448
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 17.53 6.108 2.87 0.004 5.558 29.502
goals 0 nan nan nan nan nan
Granite-3.3-8B-instruct
Pseudo R-squ Log-Likelihood
-0.0004765 -1375.9
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Llama-3.1-405B-instruct-fp8
Pseudo R-squ Log-Likelihood
0.1782 -767.54
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -64.5546 5.771 -11.186 0 -75.866 -53.243
‘custom‘ actions 95.0417 7.206 13.19 0 80.919 109.165
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 30.7633 6.989 4.402 0 17.066 44.461
goals 0 nan nan nan nan nan
Llama-3.3-70B-instruct
Pseudo R-squ Log-Likelihood
0.09968 -931.42
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -49.7385 5.343 -9.309 0 -60.211 -39.266
‘custom‘ actions 63.1362 6.384 9.89 0 50.624 75.649
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 35.2596 6.511 5.416 0 22.499 48.02
goals 0 nan nan nan nan nan
Llama-3.1-8B-instruct
Pseudo R-squ Log-Likelihood
0.003312 -1371.4
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -3.4977 2.824 -1.238 0.216 -9.034 2.038
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling 0.375 2.823 0.133 0.894 -5.157 5.907
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Mixtral-8x22B-instruct-v0.1
Pseudo R-squ Log-Likelihood
0.02272 -1062
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -23.8761 5.058 -4.72 0 -33.79 -13.962
‘custom‘ actions 50.4787 6.101 8.274 0 38.521 62.437
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 16.9506 6.157 2.753 0.006 4.883 29.018
goals 0 nan nan nan nan nan
Qwen2.5-72B-instruct
Pseudo R-squ Log-Likelihood
0.03135 -1023.5
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -22.2533 5.146 -4.324 0 -32.34 -12.167
‘custom‘ actions 56.6528 6.277 9.025 0 44.349 68.956
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 13.7726 6.279 2.193 0.028 1.465 26.08
goals 0 nan nan nan nan nan
Table A4: Logistic regression analysis for validity of plans for problems with feasible plans with concise prompt style
Deepseek-v3
Pseudo R-squ Log-Likelihood
0.1384 -723.47
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -49.3192 5.858 -8.419 0.000 -60.801 -37.837
‘custom‘ actions 94.7476 7.746 12.232 0.000 79.566 109.929
arity 0 nan nan nan nan nan
coupling 3.997e-16 3.945 1.01e-16 1.000 -7.732 7.732
slot-fillable variables 26.3420 7.395 3.562 0.000 11.848 40.836
goals 0 nan nan nan nan nan
Granite-3.3-8B-instruct
Pseudo R-squ Log-Likelihood
0.01306 -1352.2
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -5.658 2.837 -1.994 0.046 -11.219 -0.097
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -5.5151 2.842 -1.941 0.052 -11.084 0.054
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Llama-3.1-405B-instruct-fp8
Pseudo R-squ Log-Likelihood
0.1559 -713.1
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -56.0808 5.887 -9.527 0 -67.618 -44.543
‘custom‘ actions 99.9576 7.79 12.832 0 84.69 115.225
arity 0 nan nan nan nan nan
coupling 1.2189 3.931 0.31 0.757 -6.486 8.924
slot-fillable variables 25.7175 7.369 3.49 0 11.275 40.16
goals 0 nan nan nan nan nan
Llama-3.3-70B-instruct
Pseudo R-squ Log-Likelihood
-0.01675 -1300
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling 9.1485 3.029 3.02 0.003 3.211 15.086
slot-fillable variables 8.7886 3.01 2.92 0.003 2.89 14.687
goals 0 nan nan nan nan nan
Llama-3.1-8B-instruct
Pseudo R-squ Log-Likelihood
0.08154 -339.31
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -4.00E-16 17.191 -2.32E-17 1 -33.694 33.694
‘custom‘ actions -32.2708 9.824 -3.285 0.001 -51.525 -13.017
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals -99.841 20.12 -4.962 0 -139.276 -60.406
Mixtral-8x22B-instruct-v0.1
Pseudo R-squ Log-Likelihood
-0.008253 -1185.3
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 29.6337 4.892 6.057 0 20.045 39.223
goals 8.0143 4.786 1.674 0.094 -1.366 17.395
Qwen2.5-72B-instruct
Pseudo R-squ Log-Likelihood
-0.03739 -834.25
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -3.29E-15 5.972 -5.51E-16 1 -11.705 11.705
‘custom‘ actions 50.0114 7.328 6.825 0 35.648 64.374
arity 0 nan nan nan nan nan
coupling 4.4832 4.009 1.118 0.263 -3.374 12.341
slot-fillable variables 19.8769 7.324 2.714 0.007 5.522 34.232
goals 0 nan nan nan nan nan
Table A5: Logistic regression analysis for optimality of plans for problems with feasible plans with verbose prompt style
Deepseek-v3
Pseudo R-squ Log-Likelihood
0.02228 -1342.1
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -1.8931 4.346 -0.436 0.663 -10.411 6.625
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -16.0413 3.061 -5.24 0 -22.041 -10.041
slot-fillable variables 7.9159 4.59 1.725 0.085 -1.08 16.912
goals 0 nan nan nan nan nan
Granite-3.3-8B-instruct
Pseudo R-squ Log-Likelihood
0.04275 -1236.8
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions -2.7211 3.102 -0.877 0.38 -8.801 3.359
arity 0 nan nan nan nan nan
coupling -28.8463 3.278 -8.799 0 -35.272 -22.421
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Llama-3.1-405B-instruct-fp8
Pseudo R-squ Log-Likelihood
0.05037 -1236.7
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -6.761 4.742 -1.426 0.154 -16.055 2.533
‘custom‘ actions 22.0291 5.675 3.882 0 10.906 33.152
arity 0 nan nan nan nan nan
coupling -22.9507 3.236 -7.093 0 -29.292 -16.609
slot-fillable variables 21.0509 5.726 3.676 0 9.828 32.274
goals 0 nan nan nan nan nan
Llama-3.3-70B-instruct
Pseudo R-squ Log-Likelihood
0.07216 -1209.5
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -25.7043 6.722 -3.824 0 -38.88 -12.529
‘custom‘ actions 11.1087 5.649 1.966 0.049 0.037 22.181
arity 18.4486 6.866 2.687 0.007 4.992 31.905
coupling -18.039 3.248 -5.554 0 -24.405 -11.673
slot-fillable variables 28.5103 5.866 4.86 0 17.013 40.007
goals 0 nan nan nan nan nan
Llama-3.1-8B-instruct
Pseudo R-squ Log-Likelihood
0.0603 -1023.4
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions -28.2565 3.368 -8.39 0 -34.858 -21.655
arity 0 nan nan nan nan nan
coupling -30.1313 3.653 -8.248 0 -37.291 -22.971
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Mixtral-8x22B-instruct-v0.1
Pseudo R-squ Log-Likelihood
0.04701 -1302.3
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -0.3057 6.459 -0.047 0.962 -12.965 12.354
‘custom‘ actions 0 nan nan nan nan nan
arity 17.5519 6.677 2.629 0.009 4.465 30.638
coupling -21.6725 3.125 -6.935 0 -27.798 -15.547
slot-fillable variables 5.1429 4.821 1.067 0.286 -4.306 14.592
goals 0 nan nan nan nan nan
Qwen2.5-72B-instruct
Pseudo R-squ Log-Likelihood
0.05514 -1299.6
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -24.8377 3.085 -8.051 0 -30.884 -18.791
slot-fillable variables 9.6276 3.015 3.194 0.001 3.719 15.536
goals 0 nan nan nan nan nan
Table A6: Logistic regression analysis for optimality of plans for problems with feasible plans with concise prompt style
Deepseek-v3
Pseudo R-squ Log-Likelihood
0.009851 -1279.7
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 20.4360 4.515 4.527 0.000 11.588 29.284
arity 4.1622 4.290 0.970 0.332 -4.246 12.570
coupling -22.6719 3.180 -7.130 0.000 -28.904 -16.440
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Granite-3.3-8B-instruct
Pseudo R-squ Log-Likelihood
0.01306 -1352.2
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -5.658 2.837 -1.994 0.046 -11.219 -0.097
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling -5.5151 2.842 -1.941 0.052 -11.084 0.054
slot-fillable variables 0 nan nan nan nan nan
goals 0 nan nan nan nan nan
Llama-3.1-405B-instruct-fp8
Pseudo R-squ Log-Likelihood
0.1559 -713.1
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -56.0808 5.887 -9.527 0 -67.618 -44.543
‘custom‘ actions 99.9576 7.79 12.832 0 84.69 115.225
arity 0 nan nan nan nan nan
coupling 1.2189 3.931 0.31 0.757 -6.486 8.924
slot-fillable variables 25.7175 7.369 3.49 0 11.275 40.16
goals 0 nan nan nan nan nan
Llama-3.3-70B-instruct
Pseudo R-squ Log-Likelihood
-0.01675 -1300
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling 9.1485 3.029 3.02 0.003 3.211 15.086
slot-fillable variables 8.7886 3.01 2.92 0.003 2.89 14.687
goals 0 nan nan nan nan nan
Llama-3.1-8B-instruct
Pseudo R-squ Log-Likelihood
0.08154 -339.31
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -4.00E-16 17.191 -2.32E-17 1 -33.694 33.694
‘custom‘ actions -32.2708 9.824 -3.285 0.001 -51.525 -13.017
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 0 nan nan nan nan nan
goals -99.841 20.12 -4.962 0 -139.276 -60.406
Mixtral-8x22B-instruct-v0.1
Pseudo R-squ Log-Likelihood
-0.008253 -1185.3
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length 0 nan nan nan nan nan
‘custom‘ actions 0 nan nan nan nan nan
arity 0 nan nan nan nan nan
coupling 0 nan nan nan nan nan
slot-fillable variables 29.6337 4.892 6.057 0 20.045 39.223
goals 8.0143 4.786 1.674 0.094 -1.366 17.395
Qwen2.5-72B-instruct
Pseudo R-squ Log-Likelihood
-0.03739 -834.25
coef std err z $P>|z|$ [0.025 0.975]
optimal plan length -3.29E-15 5.972 -5.51E-16 1 -11.705 11.705
‘custom‘ actions 50.0114 7.328 6.825 0 35.648 64.374
arity 0 nan nan nan nan nan
coupling 4.4832 4.009 1.118 0.263 -3.374 12.341
slot-fillable variables 19.8769 7.324 2.714 0.007 5.522 34.232
goals 0 nan nan nan nan nan
LLM Plan Length Distributions
Figure A1: Distribution of LLM-generated plan lengths for different prompt styles